OpenStudy (anonymous):

solve the equation by completing the square m^2+7m-294=0

OpenStudy (anonymous):

hello can you help?

OpenStudy (anonymous):

by adding and subtracting the square of the half factor of (m) 7/2 = 3.5 (3.5)^2 = 12.25

OpenStudy (anonymous):

ya i got that but then i got confused after doing m^2+7+12.25=294+12.25

OpenStudy (anonymous):

im not understanding this thanks soo much for coming back

OpenStudy (anonymous):

(m+3.5)^2 = 306.25 so \[\sqrt{(m+3.5)^{2}} = \sqrt{306.25}\] then m+3.5 = 17.5 so m = 14 ;)

OpenStudy (anonymous):

ohhhhhhhh thankyou

OpenStudy (anonymous):

medal please ;)

OpenStudy (anonymous):

@Abdulhameed

OpenStudy (anonymous):

how do i give medals

OpenStudy (anonymous):

you`ve already gave me one lol

OpenStudy (anonymous):

ohhhh